# Explaining Explainability: Understanding Concept Activation Vectors

### v1

@@ 8 Conclusion and Future Work
In this work, we explore three key properties that influence concept activation vectors (CAVs): consistency, entanglement and spatial dependence. First, we derive conditions under which CAVs in different layers are not consistent and substantiate our findings with empirical evidence. This sheds light on why CAV-based explanations methods can give conflicting conclusions across layers. Next, we introduce a visualisation tool designed to facilitate the exploration of associations between concepts within a dataset and model. Lastly, we show that spatial dependence impacts CAVs, and introduce a method that can be used to detect spatial dependence within models. Each of these experiments was conducted using a synthetic dataset, Elements, where custom probe datasets can easily be created to analyse properties of interest. We release this dataset to help further explore this problem space.
In the introduction, we cited several interpretability methods that employ vector representations to convey semantically meaningful concepts. Our study has illuminated certain properties and consequential outcomes arising from these vector-based approaches. In future research, the characteristics inherent in alternative forms of representation, such as clusters within activation space [11], should be investigated and the relative merits assessed.
9 Acknowledgements
We appreciate both the members of OATML and the Noble group for your support and discussions during the project, in particular Andrew Jesson. We are also grateful to Been Kim for your thoughts and feedback on our work. A. Nicolson is supported by the EPSRC Centre for Doctoral Training in Health Data Science (EP/S02428X/1). J.A. Noble acknowledges EPSRC grants EP/X040186/1 and EP/T028572/1.
@@ 10 Consistency Proof
Let ${\bm{a}}_{l_{1},i}$ be the activation vector in layer $l$ for the input ${\bm{x}}_{i}\in\mathbb{X}$. Function $f$ projects the activations in layer $l_{1}$ to layer $l_{2}$, where $l_{1}<l_{2}$, i.e. $f({\bm{a}}_{l_{1},i})={\bm{a}}_{l_{2},i}$.
@@
where ${\bm{u}}\in\mathbb{R}^{m_{l_{1}}}$ and ${\bm{v}}\in\mathbb{R}^{m_{l_{2}}}$ are perturbation vectors and $m_{l_{1}}$ and $m_{l_{2}}$ are the dimensions of layer $l_{1}$ and $l_{2}$, respectively. For the two perturbations to have the same effect on the activations (and hence the model) it must hold that:
@@
Assume we have ${\bm{u}}$ and are aiming to find a ${\bm{v}}$ which satisfies Eq. 20. If $f$ conserves vector addition then
$f({\bm{a}}_{l_{1},i}+{\bm{u}})$ $=f({\bm{a}}_{l_{1},i})+{\bm{v}}$ (21)
$f({\bm{a}}_{l_{1},i})+f({\bm{u}})$ $=f({\bm{a}}_{l_{1},i})+{\bm{v}}$
${\bm{v}}$ $=f({\bm{u}})$
So, if $f$ conserves vector addition, it is possible to have consistent vectors ${\bm{v}}$ and ${\bm{u}}$ iff ${\bm{v}}=f({\bm{u}})$.
Alternatively, if $f$ does not conserve vector addition, we cannot simplify Eq. 20. In that case, we get the following solution for ${\bm{v}}$:
${\bm{v}}=f({\bm{a}}_{l_{1},i}+{\bm{u}})-f({\bm{a}}_{l_{1},i}).$ (22)
If ${\bm{v}}$ depends on $i$ then there does not exist a ${\bm{v}}$ such that Eq. 20 is true for all $i$.
10.1 ReLU Proof
@@
If $a_{l_{1},i,j}+u_{j}>0,a_{l_{1},i,j}\leq 0$ or $a_{l_{1},i,j}+u_{j}\leq 0,a_{l_{1},i,j}>0$ for any element $j$ then there does not exist a ${\bm{v}}$ such that Eq. 20 is true for all $i$, i.e., when either of these statements are true, you cannot have two vectors which have the same effect on the activations across layers for all possible inputs.
10.2 Sigmoid Proof
@@
For a single $v$ to exist which is consistent for all $a_{l_{1}}$ it cannot depend on $a_{l_{1}}$. Since the left half of Eq. 29 depends on $a_{l_{1}}$, the only way that $v$ does not depend on $a_{l_{1}}$ is if the right hand side cancels out the left. This only occurs when $u=v=0$. In other words, the only vector which is consistent across a sigmoid function is a vector which has no effect on the activations.
@@ 11 Implementation Details
In this section, we provide general implementation details applicable to the whole paper. For details relating to individual experiments and additional results, see Sections 13, 14 and 15.
@@ 11.1 Concept Activation Vectors
In [23], a statistical test, TCAV, determines whether the model’s sensitivity to a concept is significant. The test compares a set of CAV scores found using a concept dataset with CAV scores found using random data. To do this, we must find multiple CAVs for each concept. In practice, each of these CAVs is trained with the same positive set, $\mathbb{X}_{c}^{+}$, but a different random set, $\mathbb{X}_{c}^{r-}$, where $r\in{1,2\dots R}$ denotes the random index. A CAV corresponding to a specific random index is labelled ${\bm{v}}_{c,l}^{r}$.
@@
In this work, we create multiple CAVs per training run (30 unless otherwise stated), each using the same positive probe dataset but a different random set. We label a CAV trained with a specific random set as $v_{c,l}^{r}$, where $r\in{1,2\dots R}$ denotes the random index. Random CAVs are generated from pairwise combinations of random data sets, and we conduct a two-sided Welch’s t-test to test whether the means of concept and random TCAV scores are equal. If a set of CAVs passes this test with a $p$ value less than 0.010.010.010.01444we use a threshold of $0.01$ to help reduce the false discovery rate, we consider the concept meaningful. We refer to the mean TCAV score of the random CAVs as the null; it acts as the TCAV score all other CAVs should be compared against to understand their sensitivity to the model. The null is often very close to $0.5$, simplifying the interpretation of the TCAV score to the concept having positive sensitivity when greater than $0.5$ and negative when less.
@@ 11.2 Elements
The models for datasets $\mathbb{E}_{2}$ and $\mathbb{E}_{3}$ in Sec. 6.2 are the same architecture as for the simple dataset ($\mathbb{E}_{1}$).
@@ 11.3 ImageNet
We use the default weights for a ResNet-50 [20] in the TorchVision package in PyTorch, which used a variety of data augmentation techniques including Mixup [47], Cutmix [46], TrivialAugment [31], and Batch Augmentation [21].
@@
Most probe datasets used to train CAVs were collated from the Broden dataset [5], particularly focusing on textures such as striped, meshed or dotted, or objects such as car, sea or person. Some concepts were manually curated, such as the anemone concept, which was collected from test images of the ‘anemone fish’ class from ImageNet that were not used elsewhere in the experiments. Examples of some of these concepts are available in Figure 7.
@@ 11.4 Layer Selection
When creating CAVs, we need to choose a model layer. For the small CNN used for Elements, we can create CAVs for all layers, but for larger models, such as the ResNet-50 for ImageNet, it is computationally infeasible. In this paper, we focus on layers near the end of the model. The justification for this is twofold. (1) From an information theory perspective, the activations earlier in the network may contain more irrelevant information, suggesting the activations closer to the output may be more relevant to the prediction [41, 29]. We aim to use TCAV to explain the model output, therefore later layers may be more desirable. (2) The model representations may be more complex in later layers. This allows us to create CAVs for more complex concepts. We find that the empirical evidence supports these hypotheses. Figures 8 and 9 show the accuracy of the linear classifiers used to create the CAVs on a held-out test set for each probe dataset. The accuracy for each concept tends to increase in later layers. This suggests the CAVs better capture the model representations in later layers. However, we observe variation across the concepts. For example, the colours in Elements are easily classified in all layers, whereas the shapes/textures have lower performance in layers.1. As our goal is to understand the behaviour of concept vectors (when the concept is represented), we focus on CAVs that obtain at least $90\%$ test accuracy. Therefore, we omit layers.1 in our analysis.
@@ 12.2 Elements Configuration
For some experiments in Sec. 6.3 we use a spatially dependent version of the standard Elements dataset. This has the same configuration but introduces spatially dependent classes. As in the standard dataset, there is a class for all combinations of two and three concepts, e.g. ‘striped squares’ or ‘spotted cyan triangles’. However, there are additional classes which depend on where the element is present in the image. For all classes involving triangles, we add two new classes which depend on if the object is in the top or bottom half of the image. For example, the class ‘blue triangles’ will now have two additional classes related to it of ‘blue triangles on the top’ and ‘blue triangles on the bottom’. Similarly, we introduce two new classes for all classes involving squares, but for the left/right halves of the image rather than the top/bottom.
@@
Textures spots, stripes spots, stripes
@@ 13 Consistency Experiment Details
Concept CAV These are simply normal CAVs trained as described in Sections 2 and 11, i.e., the CAVs are trained using a probe dataset containing $\mathbb{X}^{+}_{c}$ and $\mathbb{X}^{-}_{c}$. The distribution is over $r_{2}$ for different random probe datasets $\mathbb{X}_{c,r_{2}}$ for ${\bm{v}}_{c,l_{2}}$, where $r_{2}\neq r_{1}$ denotes the random set.
@@ 13.1 Scaling perturbations
We performed experiments to explore the sensitivity of consistency error to the size of the perturbation, $\gamma$, for various layers and concepts for the ImageNet and Elements datasets. In Figures 12 and 13, we show the results for a variety of concepts for both the ImageNet and Elements datasets. Similar patterns were observed across experiments. As we increase $|\gamma|$, the consistency error scales linearly, as a larger perturbation causes a larger difference between $f(\hat{{\bm{a}}}_{l_{1}})$ and $\hat{{\bm{a}}}_{l_{2}}$. The scale of the y axis on the left of Figures 12 and 13 is not particularly meaningful without context, so we scale it by the norm of the perturbation in layer $l_{2}$, $||\gamma{\bm{v}}_{c,l_{2}}\frac{\overline{||{\bm{a}}_{l_{2}}||}}{||{\bm{v}}_{c,l_{2}}||}||$, on the right of the same figures. Values greater than one mean that the difference between $f(\hat{{\bm{a}}}_{l_{1}})$ and $\hat{{\bm{a}}}_{l_{2}}$ are larger than the perturbation made in layer $l_{2}$.
@@
Figure 13: The mean consistency error across 10 CAVs for striped (top), lined (middle) and dotted (bottom) CAVs across layers (left) scaled by the size of the perturbation (right) for a ResNet-50 train on ImageNet.
Figure 14: The sensitivity of the mean consistency error to scaling $\gamma$ for an optimised CAV for the penultimate convolutional layer in a ResNet50 trained on ImageNet.

### v2

@@ 8 Conclusion and Future Work
In this work, we explore three key properties that influence concept activation vectors (CAVs): consistency, entanglement and spatial dependence. First, we derive conditions under which CAVs in different layers are not consistent and substantiate our findings with empirical evidence. This sheds light on why CAV-based explanations methods can give conflicting conclusions across layers. Next, we introduce visualisations designed to facilitate the exploration of associations between concepts within a dataset and model. Lastly, we show that spatial dependence impacts CAVs, and introduce a method that can be used to detect spatial dependence within models. We provided clear recommendations on how to mitigate the impact of these properties on CAV-based explanations and demonstrated how to use those recommendations for a medical imaging use-case. The CAV properties were explored using a synthetic dataset, Elements, where custom probe datasets can easily be created to analyse properties of interest. We release this dataset to help further explore this problem space.
In the introduction, we cited several interpretability methods that employ vector representations to convey semantically meaningful concepts. Our study has illuminated certain properties and consequential outcomes arising from these vector-based approaches. In future research, the characteristics inherent in alternative forms of representation, such as clusters within activation space (Crabbé & van der Schaar, 2022), should be investigated and the relative merits assessed.
Acknowledgments
We appreciate both the members of OATML and the Noble group for your support and discussions during the project, in particular Andrew Jesson. We are also grateful to Been Kim for your thoughts and feedback on our work. Thank you to Ben Eastwood for our discussions on mathematical proofs. We also extend our thanks to the reviewers and the action editor of TMLR for their invaluable feedback and constructive comments which have enhanced the quality of the resulting paper. A. Nicolson is supported by the EPSRC Centre for Doctoral Training in Health Data Science (EP/S02428X/1). L.S. is supported by the EPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems grant (EP/S024050/1) and DeepMind. J.A. Noble acknowledges EPSRC grants EP/X040186/1 and EP/T028572/1.
@@ 9 Consistency Proof
Let ${\bm{a}}_{l,i}$ be the activation vector in layer $l$ for the input ${\bm{x}}_{i}\in\mathbb{X}$. Function $f$ maps the activations in layer $l_{1}$ to layer $l_{2}$, where $l_{1}<l_{2}$, i.e. $f({\bm{a}}_{l_{1},i})={\bm{a}}_{l_{2},i}$. We assume that $f$ is continuous and differentiable.
@@
where ${\bm{u}}\in\mathbb{R}^{m_{l_{1}}}$ and ${\bm{v}}\in\mathbb{R}^{m_{l_{2}}}$ are vectors of non-zero norm (since CAVs are directions in activation space) and $m_{l_{1}}$ and $m_{l_{2}}$ are the dimensions of layer $l_{1}$ and $l_{2}$, respectively.
For the two perturbations to have the same effect on the activations (and hence the model) it must hold that:
@@
where we have substituted in Eq. 17 and Eq. 18. For ${\bm{u}}$ and ${\bm{v}}$ to be consistent for all possible activations they must be constant with respect to ${\bm{a}}_{l_{1},i}$, i.e. we can assume that ${\bm{u}}$ and ${\bm{v}}$ are not functions of ${\bm{a}}_{l_{1},i}$. If we rearrange Eq. 19 to obtain an equation for ${\bm{v}}$, we obtain
${\bm{v}}=f({\bm{a}}+{\bm{u}})-f({\bm{a}}).$ (20)
where we have simplified the notation by writing ${\bm{a}}_{l_{1},i}$ as ${\bm{a}}$. Let us differentiate Eq. 20
$\frac{d}{d{\bm{a}}}{\bm{v}}$ $=\frac{d}{d{\bm{a}}}(f({\bm{a}}+{\bm{u}})-f({\bm{a}}))$ (21)
$\mathbf{0}$ $=\frac{d}{d{\bm{a}}}f({\bm{a}}+{\bm{u}})-\frac{d}{d{\bm{a}}}f({\bm{a}})$ (22)
$\mathbf{0}$ $=f^{\prime}({\bm{a}}+{\bm{u}})-f^{\prime}({\bm{a}})$ (23)
$f^{\prime}({\bm{a}}+{\bm{u}})$ $=f^{\prime}({\bm{a}})$ (24)
which implies that the derivative of $f$, $f^{\prime}$, is periodic with period ${\bm{u}}$. This is a strong restriction on the form that $f^{\prime}$ can take. Let’s integrate to find out what implications it has on the form of $f$. First, let’s split the periodic function $f^{\prime}$ into its mean value and its oscillatory part:
$f^{\prime}({\bm{a}})=g^{\prime}({\bm{a}})+M$ (25)
where $M\in\mathbb{R}^{m_{l_{2}}\times m_{l_{1}}}$ is the mean value across one interval for each component of $f^{\prime}({\bm{a}})$ and $g^{\prime}({\bm{a}})$ is a periodic function with zero integral across a single period, i.e.
[…]
If we integrate $g^{\prime}({\bm{a}})$ from ${\bm{a}}$ to ${\bm{a}}+{\bm{u}}$ (using a change of variables with ${\bm{t}}\in\mathbb{R}^{m_{l_{1}}}$) we find
[…]
and by using Eq. 26 we find
$g({\bm{a}}+{\bm{u}})-g({\bm{a}})=0.$ (28)
Therefore $g$ is periodic with period ${\bm{u}}$. If we now take the integral of $f^{\prime}({\bm{a}})$, we find
[…]
$f({\bm{a}})$ $=g({\bm{a}})+M{\bm{a}}+{\bm{b}}.$ (31)
Hence, $f$ satisfies Eq. 24 (and therefore Eq. 20) if and only if it is composed of a periodic function with period ${\bm{u}}$ and a linear term. However, there are further restrictions upon $f$. Let $M=\mathbf{0}$ so that $f$ is simply a periodic function. In this case
${\bm{v}}$ $=f({\bm{a}}+{\bm{u}})-f({\bm{a}})$ (32)
${\bm{v}}$ $=f({\bm{a}})-f({\bm{a}})$ (33)
${\bm{v}}$ $=\mathbf{0}$ (34)
which contradicts our non-zero assumption on the norm of ${\bm{v}}$. Hence we can obtain layer consistent vectors ${\bm{v}}$ and ${\bm{u}}$ if and only if $f$ is composed of a periodic function with period ${\bm{u}}$ and a non-zero linear term $M$. We provide no proof as to whether this form of $f$ can occur in practice in a neural network, however our empirical results in the main paper and § 12 suggest that it does not. Our proof holds generally, however, in the next three sections, we go into more detail for a linear function as it is a special case of Eq. 31 and for the ReLU and sigmoid functions as they are common activation functions used in neural networks.
9.1 Special Case: Linear Function
One counter-example that at first look seems to contradict Eq. 31 is a linear function. If $f$ is a linear function, i.e. it conserves vector addition, then Eq. 20 trivially holds:
${\bm{v}}$ $=f({\bm{a}}+{\bm{u}})-f({\bm{a}})$ (35)
${\bm{v}}$ $=f({\bm{a}})+f({\bm{u}})-f({\bm{a}})$ (36)
${\bm{v}}$ $=f({\bm{u}}).$ (37)
The general result (Eq. 31) requires that $f$ be a combination of a periodic function, $g({\bm{a}})$, and a linear function, $M{\bm{a}}+{\bm{b}}$, but we have just shown that $f=M{\bm{a}}+{\bm{b}}$ would also hold. This is because a function that outputs some constant value is a special case of a periodic function, where there is no minimal period as all periods are valid. So, for the case where $f$ is linear $g({\bm{a}})={\bm{c}}$, where ${\bm{c}}\in R^{m_{l_{2}}}$. Or for a more specific example, in the case of $f=M{\bm{a}}+{\bm{b}}$, ${\bm{c}}=\mathbf{0}$. Hence, the result in Eq. 31 generally holds and, for the case where $f$ is linear, the only layer consistent vector ${\bm{v}}$ in layer $l_{2}$ for some vector ${\bm{u}}$ in layer $l_{1}$ is that same vector projected into layer $l_{2}$ by $f$, i.e. $f({\bm{u}})$.
9.2 Example: ReLU Function
@@
If $a_{l_{1},i,j}+u_{j}>0,a_{l_{1},i,j}\leq 0$ or $a_{l_{1},i,j}+u_{j}\leq 0,a_{l_{1},i,j}>0$ for any element $j$ then there does not exist a ${\bm{v}}$ such that Eq. 19 is true for all $i$, i.e., when either of these statements are true, you cannot have two vectors which have the same effect on the activations across layers for all possible inputs. And if we assume that the elements of ${\bm{a}}$ can take any value in practice then there exists no two layer consistent vectors across a ReLU function.
9.3 Example: Sigmoid Function
@@
For a single $v$ to exist which is consistent for all $a_{l_{1}}$ it cannot depend on $a_{l_{1}}$. Since the left half of Eq. 44 depends on $a_{l_{1}}$, the only way that $v$ does not depend on $a_{l_{1}}$ is if the right hand side cancels out the left. This only occurs when ${\bm{u}}={\bm{v}}=\mathbf{0}$. Since ${\bm{u}}$ and ${\bm{v}}$ are directions in activation space, and hence have a non-zero norm, this is a contradiction. Therefore, for the sigmoid function, under no conditions does there exist layer consistent vectors.
@@ 10 Implementation Details
In this section, we provide general implementation details applicable to the whole paper. For details relating to individual experiments and additional results, see Sections 12, 13 and 14.
@@ 10.1 Concept Activation Vectors
In (Kim et al., 2018), a statistical test, TCAV, determines whether the model’s sensitivity to a concept is significant. The test compares a set of CAV scores found using a concept dataset with CAV scores found using random data. To do this, we must find multiple CAVs for each concept. In practice, each of these CAVs is trained with the same positive set, $\mathbb{X}_{c}^{+}$, but a different random set, $\mathbb{X}_{c}^{r-}$, where $r\in{1,2\dots R}$ denotes the random index. A CAV corresponding to a specific random index is labelled ${\bm{v}}_{c,l}^{r}$.
@@
In this work, we create multiple CAVs per training run (30 unless otherwise stated), each using the same positive probe dataset but a different random set. We label a CAV trained with a specific random set as $v_{c,l}^{r}$, where $r\in{1,2\dots R}$ denotes the random index. Random CAVs are generated from pairwise combinations of random data sets, and we conduct a two-sided Welch’s t-test to test whether the means of concept and random TCAV scores are equal. If a set of CAVs passes this test with a $p$ value less than 0.010.010.010.01555we use a threshold of $0.01$ to help reduce the false discovery rate, we consider the concept meaningful. We refer to the mean TCAV score of the random CAVs as the null; it acts as the TCAV score all other CAVs should be compared against to understand their sensitivity to the model. The null is often very close to $0.5$, simplifying the interpretation of the TCAV score to the concept having positive sensitivity when greater than $0.5$ and negative when less.
@@ 10.2 Elements
The models for datasets $\mathbb{E}_{2}$ and $\mathbb{E}_{3}$ in section 6.2 are the same architecture as for the simple dataset ($\mathbb{E}_{1}$).
@@ 10.3 ImageNet
We use the default weights for a ResNet-50 (He et al., 2016) in the TorchVision package in PyTorch, which used a variety of data augmentation techniques including Mixup (Zhang et al., 2018), Cutmix (Yun et al., 2019), TrivialAugment (Müller & Hutter, 2021), and Batch Augmentation (Hoffer et al., 2020).
@@
Most probe datasets used to train CAVs were collated from the Broden dataset (Bau et al., 2017), particularly focusing on textures such as striped, meshed or dotted, or objects such as car, sea or person. Some concepts were manually curated, such as the anemone concept, which was collected from test images of the ‘anemone fish’ class from ImageNet that were not used elsewhere in the experiments. Examples of some of these concepts are available in Figure 11.
@@ 10.4 Layer Selection
When creating CAVs, we need to choose a model layer. For the small CNN used for Elements, we can create CAVs for all layers, but for larger models, such as the ResNet-50 for ImageNet, it is computationally infeasible. In this paper, we focus on layers near the end of the model. The justification for this is twofold. (1) From an information theory perspective, the activations earlier in the network may contain more irrelevant information, suggesting the activations closer to the output may be more relevant to the prediction (Xu et al., 2020; McGrath et al., 2022). We aim to use TCAV to explain the model output, therefore later layers may be more desirable. (2) The model representations may be more complex in later layers. This allows us to create CAVs for more complex concepts. We find that the empirical evidence supports these hypotheses. Figures 12 and 13 show the accuracy of the linear classifiers used to create the CAVs on a held-out test set for each probe dataset. The accuracy for each concept tends to increase in later layers. This suggests the CAVs better capture the model representations in later layers. However, we observe variation across the concepts. For example, the colours in Elements are easily classified in all layers, whereas the shapes/textures have lower performance in layers.1. As our goal is to understand the behaviour of concept vectors (when the concept is represented), we focus on CAVs that obtain at least $90\%$ test accuracy. Therefore, we omit layers.1 in our analysis.
@@ 11.2 Elements Configuration
For some experiments in section 6.3 we use a spatially dependent version of the standard Elements dataset. This has the same configuration but introduces spatially dependent classes. As in the standard dataset, there is a class for all combinations of two and three concepts, e.g. ‘striped squares’ or ‘spotted cyan triangles’. However, there are additional classes which depend on where the element is present in the image. For all classes involving triangles, we add two new classes which depend on if the object is in the top or bottom half of the image. For example, the class ‘blue triangles’ will now have two additional classes related to it of ‘blue triangles on the top’ and ‘blue triangles on the bottom’. Similarly, we introduce two new classes for all classes involving squares, but for the left/right halves of the image rather than the top/bottom.
@@
Textures spots, stripes, solid spots, stripes, solid
@@ 12 Consistency Experiment Details
Concept CAV These are simply normal CAVs trained as described in Sections 2 and 10, i.e., the CAVs are trained using a probe dataset containing $\mathbb{X}^{+}_{c}$ and $\mathbb{X}^{-}_{c}$. The distribution is over $r_{2}$ for different random probe datasets $\mathbb{X}_{c,r_{2}}$ for ${\bm{v}}_{c,l_{2}}$, where $r_{2}\neq r_{1}$ denotes the random set.
@@ 12.1 Scaling perturbations
We performed experiments to explore the sensitivity of consistency error to the size of the perturbation, $\gamma$, for various layers and concepts for the ImageNet and Elements datasets. In Figures 16 and 17, we show the results for a variety of concepts for both the ImageNet and Elements datasets. Similar patterns were observed across experiments. As we increase $|\gamma|$, the consistency error scales linearly, as a larger perturbation causes a larger difference between $f(\hat{{\bm{a}}}_{l_{1}})$ and $\hat{{\bm{a}}}_{l_{2}}$. The scale of the y axis on the left of Figures 16 and 17 is not particularly meaningful without context, so we scale it by the norm of the perturbation in layer $l_{2}$, $||\gamma{\bm{v}}_{c,l_{2}}\frac{\overline{||{\bm{a}}_{l_{2}}||}}{||{\bm{v}}_{c,l_{2}}||}||$, on the right of the same figures. Values greater than one mean that the difference between $f(\hat{{\bm{a}}}_{l_{1}})$ and $\hat{{\bm{a}}}_{l_{2}}$ are larger than the perturbation made in layer $l_{2}$.
In addition to scaling $\gamma$ for both perturbations (in layer $l_{1}$ and $l_{2}$), we explored fixing the size of the perturbation in layer $l_{1}$ and varying $\gamma$ in $l_{2}$. We fixed $\gamma_{l_{1}}$ to be $0.01$ and then varied $\gamma_{l_{2}}$ from 00 to $0.02$ and measured the consistency error of the two perturbations. In Figure 18, we show these results for striped CAVs from a ResNet-50 trained on ImageNet. We repeated the experiment on two types of CAVs: (1) standard CAVs (2) optimised CAVs (as described in the previous section). In both cases having $\gamma$ fixed at $0.01$ gives a larger error, but the minimum error is not substantially different. We observe a larger difference in error between $\gamma_{l_{2}}=0.01$ and the minimum error for the standard CAVs than for the optimised CAVs. This is to be expected, as the directions of the standard CAVs are less aligned than the optimised CAVs.
@@
Figure 17: The mean consistency error across 10 CAVs for striped (top), lined (middle) and dotted (bottom) CAVs across layers (left) scaled by the size of the perturbation (right) for a ResNet-50 trained on ImageNet.
Figure 18: The sensitivity of the mean consistency error to scaling $\gamma_{l_{2}}$ for optimised CAVs (top) and standard CAVs (bottom) where $l_{1}=$layer3.0 (left) and $l_{1}=$layer4.0 (right) of a ResNet-50 trained on ImageNet. $\gamma=0.01$ is the standard perturbation made in our experiments and is marked by a grey dashed line and the $\gamma$ which gives a minimal consistency error is marked by a red dashed line.